OpenStudy (anonymous):

Evaluate the expression 10p0=

OpenStudy (anonymous):

can anyone help me out here with this I'm having trouble with it and thank you

hartnn (hartnn):

\(\huge ~_nP^r = \dfrac{n!}{(n-r)!}\implies ~_{10}P^0 = \dfrac{10!}{(10-0)!}\\ \huge =\dfrac{10!}{10!}=1 \)

hartnn (hartnn):

@pugsy_joy